OpenStudy (anonymous):

A curve passes through the point (0, 8) and has the property that the slope of the curve at every point P is t?

OpenStudy (anonymous):

as the slope is constant, graph is a straight line y=tx+c y=8 while x=0 so, 8=c so, y=tx+8

OpenStudy (anonymous):

@antwan , u got it?

OpenStudy (kropot72):

An equation that meets the requirements is \[y=e ^{x}+7\] and \[t=e ^{x}\] \[f \prime y=e ^{x}\]

OpenStudy (anonymous):

so let that curve be the function y(t). So we also know that dy/dt=t dy=tdt integrating will give us y+k=t^2/2+c y=t^2/2+c-k c-k is just also a constant so we can write that as c' which is a constant y=t^2/2+c' now we need to find the actual solution by finding c' by applying the initial condition y(0)=8, 8=(0)^2/2+c' c'=8 so our curve will be y=t^2/2+8

OpenStudy (anonymous):

its a Differential equation problem

OpenStudy (anonymous):

if t is constant, the curve is line. so: y-0=t(x-8) is the equation